The following OT tableau was created in L^{A}TEX. Note that the column widths and row heights are chosen automatically.

If you change the star patterns or reorder the columns, you can pipe just the 6 candidate lines through the ”shade” script to automatically revise the exclamation points, shading, and pointy hand.

(syll. troch.)  NonUnary Spread ShortFt
candidate 1 *!
candidate 2  *
candidate 3  *!  *
giant *!  *
candidate 5  *!*  **
candidate 6  *

Warning: This document should print correctly. But if you are viewing it with xdvi, then the shaded cells may appear as solid black. If you are viewing it with ghostview or Acrobat Reader (PDF), then the shading may appear to obscure some of the gridlines unless you zoom in a lot.

Warning: The pointy hand should appear correctly in PDF. However, in the other formats (.ps and .dvi) it may look wrong if you don’t have the necessary font installed.